### v1

# PBLean: Pseudo-Boolean Proof Certificates for Lean 4

Abstract
We present PBLean, a method for importing VeriPB pseudo-Boolean (PB) proof certificates into Lean 4. Key to our approach is reflection: a Boolean checker function whose soundness is fully proved in Lean and executed as compiled native code. Our method scales to proofs with tens of thousands of steps that would exhaust memory under explicit proof-term construction. Our checker supports all VeriPB kernel rules, including cutting-plane derivations and proof-by-contradiction subproofs. In contrast to external verified checkers that produce verdicts, our integration yields Lean theorems that can serve as composable lemmas in larger formal developments. To derive theorems about the original combinatorial problems rather than about PB constraints alone, we support verified encodings. This closes the trust gap between solver output and problem semantics since the constraint translation and its correctness proof are both formalized in Lean. We demonstrate the approach on various combinatorial problems.
1 Introduction
Pseudo-Boolean (PB) solvers handle linear constraints over Boolean variables and have become prominent tools for combinatorial optimization and verification [11]. Since these solvers carry out sophisticated derivations, independent verification of their results is particularly desirable. However, PB reasoning involves cutting-planes proof steps that resolution-based formats cannot capture efficiently.
Haken’s classical result [15] shows that the pigeonhole principle requires exponential-size resolution proofs, whereas polynomial-size cutting-plane proofs exist [6]. Beyond pigeonhole, cutting planes also efficiently capture XOR/parity constraints [14] and certified symmetry breaking [4], which are reasoning methods that lack efficient resolution-based proof logging. This motivated the VeriPB proof format [10, 4], based on cutting planes and now used by solvers such as RoundingSat for certified PB solving.
The LRAT format [16] enables SAT solvers to produce clausal certificates that can be checked in theorem provers. Lean’s FromLRAT module [23] imports such certificates and produces kernel-checked theorems. Heule and Scheucher [17] established the empty hexagon theorem using SAT solving with verified certificate checking; Subercaseaux et al. [24] subsequently formalized this encoding in Lean and verified the reduction from geometry to SAT. A similar integration for pseudo-Boolean reasoning has been missing.
Why not compile PB reasoning to clausal proofs and reuse LRAT? Bryant et al. [5] showed that this is possible via BDD compilation, and their approach handles classical hard cases such as pigeonhole and parity efficiently. However, the compilation requires coordinating a PB solver with a BDD-based clausal proof generator, and whether it preserves efficiency for all PB reasoning remains open. Direct VeriPB import avoids this detour.
The CakePB project [13] provides a verified checker extracted from HOL4 via CakeML, producing a standalone binary that outputs verdicts. The main advantage of CakePB is its small, trusted computing base enabled by verified compilation. However, the verified result (a binary verdict) cannot serve as a lemma in larger formal developments.
A proof certificate establishes that certain PB constraints are unsatisfiable, but says nothing about whether those constraints correctly model the intended problem. An encoding bug can produce unsatisfiable constraints even when solutions exist. We address this gap with verified encodings: both the constraint translation and its correctness proof are entirely in Lean, so that certificate import yields theorems about the original problem, not merely about PB constraints.
We present PBLean, a VeriPB integration for Lean 4 that parses kernel-format proofs and verifies them via reflection. Instead of building Lean proof terms for each derivation step, we define a Boolean checker function and prove its soundness once; Lean’s native_decide then runs the checker as compiled code. This follows the pattern of Lean’s bv_decide tactic [21] and scales to instances that would exhaust memory under explicit proof construction. Our implementation covers all VeriPB kernel rules, including proof-by-contradiction subproofs for optimization reasoning. We apply the verified encoding workflow to independence numbers of Paley graphs and other combinatorial theorems to obtain end-to-end-verified results.
2 Background
A pseudo-Boolean (PB) constraint has the form $\sum_{i=1}^{n}a_{i}\ell_{i}\geq d$, where $a_{i},d\in\mathbb{N}$ are non-negative coefficients, and each $\ell_{i}$ is a literal (a variable $x_{i}$ or its negation $\overline{x}_{i}$). Under an assignment $\sigma$, a literal evaluates to 1 if satisfied and 0 otherwise. The constraint is satisfied when the weighted sum reaches the degree $d$. For example, $3x_{1}+5x_{2}\geq 4$ is satisfied by $\sigma(x_{1})=\sigma(x_{2})=1$ (sum $8\geq 4$) but not by $\sigma(x_{1})=1$, $\sigma(x_{2})=0$ (sum =3<4=3<4). Clauses are a special case: $x_{1}\vee x_{2}\vee\overline{x}_{3}$ corresponds to $x_{1}+x_{2}+\overline{x}_{3}\geq 1$.
Cutting-planes reasoning derives new constraints via: (1) addition of constraints, (2) scalar multiplication by $k\in\mathbb{N}^{+}$, (3) division by $k$ with the degree rounded up, and (4) saturation, which caps each coefficient at the degree. These rules strictly generalize resolution [6]: resolution corresponds to addition with complementary literal cancellation. The power of cutting planes stems from the interplay of multiplication and division, which derives constraints unreachable by resolution alone [25].
The VeriPB proof format [10, 4] provides two modes: an augmented mode with high-level rules (symmetry breaking, redundance-based strengthening) and a kernel mode (version 3.0) with only primitive operations. The kernel format’s small rule set of just six rule types enables a rigorous soundness proof and makes it well-suited for theorem prover integration. It supports: pol (polynomial operations including all cutting-planes rules), rup (reverse unit propagation with conflict hints), pbc (proof by contradiction with inline subproofs), del (constraint deletion, sound by monotonicity), sol/soli (solution logging), and conclusion (UNSAT/SAT/BOUNDS verdict). Solvers such as RoundingSat produce augmented proofs; veripb --elaborate translates them to kernel format.
Lean 4 [7] provides a metaprogramming framework where tactics run in monads such as MetaM. Code in MetaM constructs Expr values (Lean’s internal representation of proof terms), and, upon type-checking, Lean’s kernel verifies their correctness. Our approach builds Expr values from VeriPB proofs; the parser and proof-term constructor are untrusted code, as in FromLRAT. The trusted base consists of Lean’s kernel and the soundness lemmas presented in Section 3.
3 Lean Formalization
For constraint representation, we use natural number arithmetic throughout. A $\mathit{Literal}$ is $\mathtt{pos}\ i$ or $\mathtt{neg}\ i$, and a $\mathit{Term}$ is a pair $(a,\ell)\in\mathbb{N}\times\mathit{Literal}$; a $\mathit{Constr}$ is a term list plus degree. Given a valuation $v:\mathbb{N}\to\mathit{Bool}$, we define $\mathit{evalLit}(v,\ell)\in\{0,1\}$ (1 if satisfied) and $\mathit{evalSum}(v,T)=\sum_{(a,\ell)\in T}a\cdot\mathit{evalLit}(v,\ell)$ recursively to support structural induction; $\mathit{coeffSum}(T)=\sum_{(a,\ell)\in T}a$ is the sum of coefficients. A constraint $(T,d)$ is satisfied when $d\leq\mathit{evalSum}(v,T)$.
Our 13 soundness lemmas are summarized in Table 1; every lemma is proved in Lean and verified by the kernel. Two details matter for matching VeriPB’s semantics: normalization must precede saturation and division, and the negation of $\sum a_{i}\ell_{i}\geq d$ is $\sum a_{i}\bar{\ell}_{i}\geq(\sum a_{i})-d+1$.
Table 1: Soundness lemmas for VeriPB kernel operations.
Operation Lean Type
Cutting planes
add_sat $c_{1}.\mathit{sat}\ v\to c_{2}.\mathit{sat}\ v\to(c_{1}.\mathit{terms}\mathbin{++}c_{2}.\mathit{terms},\ c_{1}.\mathit{deg}+c_{2}.\mathit{deg}).\mathit{sat}\ v$
mul_sat $c.\mathit{sat}\ v\to(\mathit{map}\ (\lambda(a,\ell).\ (k\cdot a,\ell))\ c.\mathit{terms},\ k\cdot c.\mathit{deg}).\mathit{sat}\ v$
div_sat $0<k\to c.\mathit{sat}\ v\to(\mathit{map}\ (\lambda(a,\ell).\ (\lceil a/k\rceil,\ell))\ c.\mathit{terms},\ \lceil c.\mathit{deg}/k\rceil).\mathit{sat}\ v$
saturate_sat $c.\mathit{sat}\ v\to(\mathit{map}\ (\lambda(a,\ell).\ (\min(a,c.\mathit{deg}),\ell))\ c.\mathit{terms},\ c.\mathit{deg}).\mathit{sat}\ v$
Normalization
cancel_pair_sat $\min(a,b)\leq d\to(\ldots(a,\ell)\ldots(b,\bar{\ell})\ldots,\ d).\mathit{sat}\ v\to(\ldots(a{-}m,\ell)\ldots(b{-}m,\bar{\ell})\ldots,\ d{-}m).\mathit{sat}\ v$
remove_zero_sat $(\mathit{pre}\mathbin{++}(0,\ell){::}\mathit{post},\ d).\mathit{sat}\ v\to(\mathit{pre}\mathbin{++}\mathit{post},\ d).\mathit{sat}\ v$
merge_terms_sat $(\ldots(a,\ell)\ldots(b,\ell)\ldots,\ d).\mathit{sat}\ v\to(\ldots(a{+}b,\ell)\ldots,\ d).\mathit{sat}\ v$
Axioms and contradiction
lit_axiom_pos/neg $\forall v.\ ([(1,\ell)],\ 0).\mathit{sat}\ v$
contra_unsat $\mathit{coeffSum}(c)<c.\mathit{deg}\to\neg\ c.\mathit{sat}\ v$
Negation and proof by contradiction
negate_sat_of_not $c.\mathit{deg}\leq\mathit{coeffSum}(c)\to\neg\ c.\mathit{sat}\ v\to c.\mathit{negate}.\mathit{sat}\ v$
propagation_forces $(T,d).\mathit{sat}\ v\to\mathit{coeffSum}(T\setminus\{(a,\ell)\})<d\to\mathit{evalLit}(v,\ell)=1$
pbc_sound $(\forall v.\ F.\mathit{allSat}\ v\to C.\mathit{negate}.\mathit{sat}\ v\to\bot)\to F.\mathit{proof}\ C$
We give two verification approaches with complementary trade-offs.
Explicit proof terms.
The first approach builds Lean Expr terms by composing soundness lemmas. For pol steps, each token in the reverse Polish notation (RPN) expression produces an application: ID lookups retrieve earlier proofs, addition applies add_sat, multiplication applies mul_sat, and so on. For rup steps, we negate the target, run PB unit propagation to find a conflicting hint, and then build a proof accumulator. For pbc subproofs, we extend the context with the negated assumption and verify inner steps recursively. The resulting Expr is then type-checked by Lean’s kernel, producing a theorem. Its trusted base consists of just the kernel and the soundness lemmas. This approach has the disadvantage that the accumulated proof terms exhaust memory for instances with thousands of constraints.
Reflection.
To overcome this limitation, our second approach follows the pattern of Lean’s bv_decide [21]. We define a Boolean function $\mathit{check}(F,\pi)$ that takes a constraint set $F$ and a proof string $\pi$, replays all derivation steps, and returns $\mathit{true}$ iff a contradictory constraint is derived. We then prove a soundness theorem: if $\mathit{check}(F,\pi)=\mathit{true}$ then $F$ is unsatisfiable. This theorem is fully proved in Lean without custom axioms. Lean’s native_decide runs the checker as compiled code, yielding a kernel proof term consisting solely of the soundness theorem applied to the computation witness. The resulting theorem has the same trust basis as bv_decide and omega—namely, Lean’s standard axioms (propext, Classical.choice, Quot.sound) plus Lean.trustCompiler, and scales to proofs with thousands of steps.
4 Implementation
PBLean is a standalone Lean 4 project (v4.28.0) with no Mathlib dependency. The kernel layer (PseudoBoolean.lean, $\scriptstyle\sim$700 lines) consists of types and 13 soundness lemmas. The code is available on GitHub. [^1] https://github.com/leansolving/pblean
The reflection-based checker (Reflect.lean, $\scriptstyle\sim$1800 lines) uses a line-by-line tokenizer to convert VeriPB proofs into an abstract syntax tree, handling nested pbc subproofs recursively. It maintains a constraint database organized as a hash map, with an @[implemented_by] annotation providing an array-based runtime replacement for performance without weakening the proof. Each rule updates or queries this database: pol evaluates the RPN expression, rup derives a conflict by negating the target and applying PB unit propagation, and pbc saves the database, adds the negated assumption, verifies inner steps recursively, then restores the snapshot. The checker returns true only when a contradictory constraint arises, and we establish soundness by induction over proof steps, connecting the Boolean computation to the kernel lemmas.
To exploit the integration, one provides an OPB file and a kernel-format proof file. The command veripb_reflect parses both, runs the checker via native_decide, and registers the resulting unsatisfiability theorem under a given name. For domain-specific applications, wrapper commands such as independent_set_reflect additionally generate the encoding from a problem instance and compose the bridge theorem automatically.
We demonstrate our approach on a problem where the encoding is also formalized in Lean, closing a trust gap. For SAT, encoding correctness was only recently formalized [24]; for PB reasoning, no such formalization existed. Figure 1 shows the overall workflow.
ProblemInstanceTrustedEncodingPB Solver+ VeriPBReflectionCheckerLeanTheoremOPBKernel Proofbridgeuntrusted
Figure 1: End-to-end workflow for trusted encodings. OPB (Optimized Pseudo-Boolean) is the standard input format for PB solvers. Solid boxes are verified in Lean; the dashed box is untrusted. The dotted arrow represents the bridge theorem composing encoding correctness with the PB refutation.
A verified encoding consists of three components: (1) an encode function translating problem instances to PB constraints, (2) a soundness theorem proving that unsatisfiability of the encoding implies non-existence of solutions, and (3) a bridge theorem composing encoding correctness with the imported VeriPB certificate to yield a theorem about the original problem.
As an example, we consider the independence number (the size of a largest independent set) of Paley graphs. For a prime $p\equiv 1\pmod{4}$, the Paley graph $\mathrm{Paley}(p)$ is defined as the graph with the vertex set $\mathbb{Z}_{p}$ and edges $(i,j)$ whenever jij-i is a nonzero quadratic residue. These graphs are self-complementary with an independence number $\alpha$ of at most $\sqrt{p}$ by the Hoffman eigenvalue bound. The encoding for “does $G$ have an independent set of size $\geq k$?” uses
$\forall(u,v)\in E(G)$ $:\bar{x}_{u}+\bar{x}_{v}\geq 1\quad\text{(independence)}$
$:x_{1}+\cdots+x_{n}\geq k\quad\text{(cardinality)}$
The Lean formalization establishes that if these constraints are unsatisfiable, then $G$ has no independent set of size $\geq k$. Combining the VeriPB certificate with this result yields a theorem about the graph. The independent_set_decide command automates the whole pipeline: given a graph and a target size $k$, it produces a Lean theorem of the form $\lnot\mathit{hasIndependentSet}(G,k)$.
5 Evaluation
We evaluate PBLean by measuring wall-clock time on an Apple M2 with cached .olean files. Measured times include parsing, checker execution via native_decide, and kernel verification. We use RoundingSat for solving and veripb --elaborate to produce kernel-format proofs.
Table 2 shows results for independence numbers of $\mathrm{Paley}(p)$ for primes $p\equiv 1\pmod{4}$ from 13 to 101.
The proofs rely on cutting-planes (pol) and reverse unit propagation (rup) steps; the largest instance (2525 edges, about 63 000 proof lines) verifies in about 3.5 minutes.
Table 2: Independent set on Paley graphs. $p$: prime, $\alpha$: independence number. Vars: PB variables, Cstrs: PB constraints. Direct: explicit proof-term construction. Reflect: reflection-based checker. TO = timeout at 60 000 msms.
$p$ $\alpha$ Vars Cstrs Solve ($ms$) Elab ($ms$) Lines Direct ($ms$) Reflect ($ms$)
13 3 13 40 59 12 35 1 150 802
17 3 17 69 38 4 66 2 310 802
29 4 29 204 53 5 233 19 922 926
37 4 37 334 105 6 554 TO 1 314
41 5 41 411 131 6 704 TO 1 536
53 5 53 690 421 13 1 848 TO 3 494
61 5 61 916 850 24 4 151 TO 6 743
73 5 73 1 315 3 145 64 9 808 TO 20 616
89 5 89 1 959 17 166 234 36 795 TO 96 234
97 6 97 2 329 21 673 272 43 214 TO 115 197
101 5 101 2 526 29 267 385 62 924 TO 200 460
As an unexpected discovery worth noting, the data yields a formally verified proof that $\alpha(\mathrm{Paley}(p))$ is not monotone in $p$: $\alpha(\mathrm{Paley}(97))=6>5=\alpha(\mathrm{Paley}(101))$. Here, each equality combines the VeriPB upper bound with a lower bound from an explicit witness verified by native_decide.
theorem paley_alpha_not_monotone :
$\exists$$ p q : Nat, p < q $\wedge$$ $\exists$$ a b : Nat,
independenceNumber (paley p) a $\wedge$$
independenceNumber (paley q) b $\wedge$$ a > b
As noted in Section 3, the explicit proof-term approach does not scale: the Direct column shows that $\mathrm{Paley}(29)$ already takes 20 s, and all instances with $p\geq 37$ exceed the 60 s timeout. In contrast, the reflection-based checker verifies $\mathrm{Paley}(53)$ in 3.5 s and $\mathrm{Paley}(101)$ in 200 s. The bottleneck appears to be the compiled checker execution rather than Lean’s kernel, which only checks the soundness theorem application.
Beyond independence numbers of Paley graphs, we provide trusted encodings for five additional combinatorial problems. These applications demonstrate the generality of the verified encoding approach. The underlying theorems are well known; the contribution lies in the end-to-end formal verification within Lean. Table 3 reports the same pipeline metrics as Table 2.
Table 3: Additional verified combinatorial results, using the reflection-based checker. Column headers as in Table 2.
Problem Vars Cstrs Solve ($ms$) Elab ($ms$) Lines Reflect ($ms$)
Langford $\neg L(2,6)$ 72 463 152 12 223 1 221
Schur $S(2)=4$ 6 12 35 4 17 843
VdW $W(2,3)=9$ 10 32 37 4 19 867
Ramsey $R(3,3)=6$ 15 40 37 4 31 894
Eq. col. $\chi_{eq}(K_{3,3,1})\!=\!5$ 28 117 72 5 490 957
Langford pairing.
A Langford pairing of order $n$ places each integer $k\in\{1,\ldots,n\}$ into a sequence of length $2n$ such that the two occurrences of $k$ are exactly $k+1$ positions apart. Langford pairings exist if and only if $n\equiv 0$ or $3\pmod{4}$.
[…]
The impossibility of $L(2,6)$ is proved via the reflection checker from a 223-line kernel proof (72 variables, 463 constraints).
Schur number.
A 2-coloring of $\{1,\ldots,n\}$ is Schur-free if no monochromatic triple (a,b,a+b)(a,b,a{+}b) exists. The Schur number $S(2)$ is the largest $n$ for which there exists such a coloring.
theorem schur_number_2 : schurNumber 4
The lower bound uses the witness coloring $\{1,4\}\mapsto\mathrm{true}$, $\{2,3\}\mapsto\mathrm{false}$. The upper bound ($\{1,\ldots,5\}$ has no Schur-free coloring) is verified via the reflection checker from a 13-line kernel proof.
Van der Waerden number.
A 2-coloring of $\{1,\ldots,n\}$ is AP-free if no monochromatic 3-term arithmetic progression (a,a+d,a+2d)(a,a{+}d,a{+}2d) exists. The Van der Waerden number $W(2,3)$ is the largest $n$ for which there exists such a coloring.
theorem vdw_number_2_3 : vanDerWaerdenNumber 8
The lower bound uses the witness RRBBRRBB; the upper bound ($\{1,\ldots,9\}$ has no AP-free coloring) is verified via the reflection checker in a 19-line kernel proof.
Ramsey number.
A 2-coloring of the edges of the complete graph $K_{n}$ is Ramsey-free if it contains no monochromatic triangle. The Ramsey number $R(3,3)$ is the smallest $n$ such that every 2-coloring of $K_{n}$ contains a monochromatic triangle.
theorem ramsey_3_3 : ramseyNumber 5
The lower bound uses the pentagon ($C_{5}$) coloring of $K_{5}$. The upper bound—that $K_{6}$ admits no Ramsey-free coloring—is verified via the reflection checker from a 31-line kernel proof derived from 15 edge variables and 40 constraints (20 triangles).
Equitable chromatic number.
A proper $k$-coloring of a graph $G$ is equitable if all color classes have size in $\{\lfloor n/k\rfloor,\lceil n/k\rceil\}$. The equitable chromatic number $\chi_{eq}(G)$ is the smallest such $k$. The balance constraints are natively pseudo-Boolean: they express cardinality bounds as linear inequalities, which would require auxiliary encoding variables in pure SAT.
theorem k331_eq_chromatic : equitableChromaticNumber k331 5
The complete tripartite graph $K_{3,3,1}$ (7 vertices, 15 edges) has standard chromatic number $\chi=3$ but equitable chromatic number $\chi_{eq}=5$, demonstrating a gap of 2. The upper bound uses an explicit 5-coloring witness; the lower bound ($\neg\mathit{hasEquitableColoring}\ K_{3,3,1}\ 4$) is verified via the reflection checker from a 490-line kernel proof with 96 proof-by-contradiction subproofs.
6 Related Work
CakePB [13] provides end-to-end verified subgraph solving in HOL4/CakeML, including both a PB proof checker and verified graph-problem encoders compiled to a standalone binary. As discussed in Section 1, the approaches are complementary: CakePB targets standalone verification with a small trusted computing base, while PBLean yields Lean theorems that can serve as lemmas in larger formal developments. Several extensions of the VeriPB framework are known, certifying MaxSAT preprocessing [18], solution-improving search [1], dynamic programming and decision diagrams [8], and pseudo-Boolean optimization [19].
An alternative to native PB proof checking is to compile PB reasoning to clausal proofs. Bryant et al. [5] showed that this is possible via BDD conversion, and their approach efficiently handles pigeonhole and parity instances via extension variables. Direct VeriPB import provides a simpler pipeline that works with the native proof format of PB solvers.
Approaches for SAT certificate checking in theorem provers include Heule et al.’s verified LRAT checker in ACL2 [16], Lammich’s GRAT in Isabelle/HOL [20], and Lean’s FromLRAT [23], whose architecture our integration closely follows. Heule and Scheucher [17] established the empty hexagon theorem via SAT solving with verified proof checking; Subercaseaux et al. [24] formalized the encoding in Lean and verified the reduction from geometry to SAT.
Rather than checking proof certificates, one can verify the solver itself. Blanchette et al. [3] formalized CDCL in Isabelle/HOL as a chain of refinements from abstract calculus to executable code; Fleury and Lammich [12] extended this to IsaSAT, a verified SAT solver. SMTCoq [9] imports SMT witnesses into Coq. Besson’s micromega [2] uses reflexive certificate checking for linear arithmetic in Coq, based on Farkas’ Lemma rather than cutting planes, and does not import external PB proofs. A related project [22] formalizes general CSP definitions in Lean 4 with verified translations to MiniZinc and SMT-LIB solver formats. PBLean extends verified PB checking to Lean 4, complementing existing SAT checking in Isabelle and PB checking in HOL4.
7 Conclusion
We propose a Lean 4 integration for VeriPB pseudo-Boolean proof certificates. The key to PBLean’s scalability is reflection: a Boolean checker with a fully proved soundness theorem, evaluated natively by the kernel. As demonstrated in Section 5, this verifies proofs with tens of thousands of steps that would exhaust memory under explicit proof construction. Our largest instance, $\mathrm{Paley}(101)$ with about 63 000 proof lines, verifies in about 3.5 minutes.
With trusted encodings, users obtain end-to-end verified combinatorial results: the independence number of $\mathrm{Paley}(101)$ is a Lean proposition, not an external verdict. This composability, not directly available with standalone verified checkers, enables PB reasoning to serve as a building block in larger formalizations.
Future work includes a veripb tactic for proof automation and integration with other Lean formalizations.
Acknowledgments
Research was carried out within the Cluster of Excellence Bilateral AI (10.55776/COE12) of the Austrian Science Funds (FWF).